### v1

@@ 1 Introduction
Nonetheless, the early exercise feature of American options poses significant challenges for computing price and Greeks, especially in high dimensions. One of the most popular methods for high-dimensional American option pricing is the least squares Monte Carlo (LSM) method [15, 22]. Computing Greeks in high dimensions is more involved, and further developments with LSM have been proposed [24, 5]. In this work, building on LSM, we shall develop a simple, fast, and accurate algorithm, termed as gradient-enhanced least squares Monte Carlo (G-LSM) method, c.f. Algorithm 1, for computing price and Greeks simultaneously at all time steps for dimensions up to 100. The key methodological innovations includes using sparse Hermite polynomial space with a hyperbolic cross index set as the ansatz space for approximating the continuation value functions (CVFs), and incorporating the gradient information for computing the expansion coefficients.
We elaborate on the two methodological innovations. First, the main obstacle of using gradient information lies in the computational expense: a $d$-variate price has $d$ partial derivatives, which grows quickly with $d$, especially when the derivative evaluation is costly. With the proposed sparse Hermite polynomial ansatz space, the derivatives of polynomial bases can be obtained at almost no extra cost, cf. Eq. 9 below. This allows greatly reducing the computational cost. Second, although using a polynomial ansatz space for the CVFs as LSM, G-LSM constructs the expansion coefficients via solving a linear least squares problem enhanced by the gradient (and hence the name G-LSM), by minimizing the mean squared error between the approximate and exact value functions at $t_{k+1}$, c.f. Eq. 15 below. This differs markedly from LSM, which approximates the conditional expectations by projection and minimizes the mean squared error of approximating CVF at $t_{k}$. Numerical experiments show that this choice can achieve better accuracy in price, Greeks, and optimal exercise strategies than LSM.
@@
Throughout, bold and plain letters represent multi-variable and scalars, respectively, and the capital and bold letters $\mathbf{S}$, $\tilde{\mathbf{X}}$ and $\mathbf{W}$ denote random vectors. The notation $\mathbf{a}\cdot\mathbf{b}$ denotes the dot product of two vectors $\mathbf{a}$ and $\mathbf{b}$, $\operatorname{Tr}(A)$ the trace of a square matrix $A$, and vector transpose. The notation $\nabla_{\mathbf{x}}f(t,\mathbf{x})$ and $\operatorname{Hess}_{\mathbf{x}}f(t,\mathbf{x})$ denote respectively the gradient and Hessian of $f$ with respect to $\mathbf{x}$. For multi-index $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{d})^{\top}\in\mathbb{N}_{0}^{d}$, $|\bm{\alpha}|=\alpha_{1}+\dots+\alpha_{d}$. For a positive-valued integrable function $\omega:\mathbb{R}^{d}\to\mathbb{R}^{+}$, the weighted space $L^{2}_{\omega}(\mathbb{R}^{d})$ is defined by
@@ 2.1 Option pricing and Greeks
Numerically, the price of Bermudan option is used to approximate the American one. The Bermudan option can be exercised at finite discrete times $0=t_{0}<t_{1}<\dots<t_{N}=T$ with $\Delta t:=t_{k+1}-t_{k}$ for all $k=0,1,\dots,N-1$. Using dynamic programming principle or Snell envelope theory [20, Section 1.8.4], the Bermudan price function ${v}_{t_{k}}$ at time $t_{k}$ is given by the following backward induction:
@@
$v_{t_{k}}(\mathbf{s})$ =$=\begin{cases}g_{t_{k}}(\mathbf{s}),\quad\text{ if }g_{t_{k}}(\mathbf{s})\geq\tilde{c}_{t_{k}}(\mathbf{s}),\\ \tilde{c}_{t_{k}}(\mathbf{s}),\quad\text{ if }g_{t_{k}}(\mathbf{s})<\tilde{c}_{t_{k}}(\mathbf{s}),\end{cases}\quad\text{ for }k=N-1:-1:0,$
@@
The options delta and gamma are defined to be the first and second order derivatives of the price function $v_{t_{k}}$ with respect to the price of underlying assets, i.e.,
(3) $\Delta_{t_{k}}$ $:=\nabla v_{t_{k}}(\mathbf{s})=\left(\frac{\partial v_{t_{k}}(\mathbf{s})}{\partial s_{1}},\dots,\frac{\partial v_{t_{k}}(\mathbf{s})}{\partial s_{d}}\right)^{\top}\quad\mbox{and}\quad\Gamma^{ij}_{t_{k}}$ $:=\frac{\partial^{2}v_{t_{k}}(\mathbf{s})}{\partial s_{j}\partial s_{i}}.$
We consider all exercise and continuation values of Bermudan option discounted to the time $t=0$.
@@ 3 Sparse Hermite polynomial expansion and gradient
    Let $\omega_{t}$, $t>0$, be the Gaussian density function defined by $\omega_{t}(\mathbf{y}):=\prod_{j=1}^{d}\rho_{t}(y_{j})$ with $\rho_{t}(y):=\frac{1}{\sqrt{2\pi t}}\exp(-\frac{y^{2}}{2t})$. The CVF $c_{t_{k}}$ satisfies $c_{t_{k}}\in{L}^{2}_{\omega_{t_{k}}}(\mathbb{R}^{d})$, since the payoff $g(\mathbf{S}_{t_{k}})\in L^{2}(\Omega,\mathcal{F}_{t_{k}},\mathbb{Q})$ has a finite second-order moment. Thus, as an orthonormal basis for $L^{2}_{\omega_{t_{k}}}(\mathbb{R}^{d})$, the set of normalized and generalized Hermite polynomials emerges as a natural choice.
@@
    The CVF $c_{t_{k}}\in{H}^{m}_{\omega_{t_{k}}}(\mathbb{R}^{d})$ for any positive integer $m$ by [25, Lemmas 4.2 and 4.3]. The smoothness of the function implies efficient polynomial approximation.
Next, we recall one-dimensional normalized and generalized Hermite polynomials $H_{n}^{(t)}(y)$ and the tensorized $d$-dimensional polynomials $H_{\bm{\alpha}}^{(t)}(\mathbf{y})$, $\bm{\alpha}\in\mathbb{N}_{0}^{d}$. For $d=1$, using the standard Hermite polynomials $H_{n}$, we define the $n$th-order normalized and generalized Hermite polynomial $H_{n}^{(t)}$ by
@@ 4.1 Gradient-enhanced Least Squares (G-LS)
First we derive a linear backward stochastic differential equation (BSDE) for the CVF $c_{k}$.
@@
This is a direct consequence of martingale representation theorem [26, Theorem 2.5.2]. For the sake of completeness, we provide a brief proof. By the Feynman-Kac formula, $f(t,\mathbf{w})$ satisfies a $d$-dimensional parabolic PDE subject to a terminal condition
[…]
@@
Next, we construct an approximation of the CVF ${c}_{k}$ by matching the terminal condition over the interval $[t_{k},t_{k+1}]$. By Theorem 4.1, the terminal value over $[t_{k},t_{k+1}]$ can be approximated by the Euler discretization, i.e.,
@@ 4.2 Gradient-enhanced Least Squares Monte Carlo (G-LSM)
where $\{\mathbf{W}^{m}_{k}\}_{m=1}^{M}$ are $M$ independent paths of the Gaussian random process $\mathbf{W}_{k}$ and $\Delta\mathbf{W}^{m}_{k}:=\mathbf{W}^{m}_{k+1}-\mathbf{W}^{m}_{k}$ is the $m$th path of increment $\Delta\mathbf{W}_{k}$. Let $A_{k}\in\mathbb{R}^{M\times N_{b}}$, $\bm{\beta}_{k}\in\mathbb{R}^{N_{b}}$ and $\mathbf{\hat{u}}_{k+1}\in\mathbb{R}^{M}$ with their components defined by
$(A_{k})_{mn}$ $=\phi_{n}^{k}(\mathbf{W}^{m}_{k})+\nabla\phi_{n}^{k}(\mathbf{W}^{m}_{k})\cdot\Delta\mathbf{W}^{m}_{k},$
@@
$\bm{\beta}_{k}=\operatorname*{argmin}_{\bm{\beta}}\|A_{k}\bm{\beta}-\mathbf{\hat{u}}_{k+1}\|_{2}^{2}=(A_{k}^{\top}A_{k})^{-1}A_{k}^{\top}\mathbf{\hat{u}}_{k+1}.$
@@
9:     Update $(\mathbf{\hat{u}}_{k})_{m}=\begin{cases}e^{-rk\Delta t}g(\mathbf{S}^{m}_{k}),&\text{if exercise}\\ (\Phi_{k}\bm{\beta}_{k})_{m},&\text{if continue}\end{cases}\text{ and }\tau_{\star}=\begin{cases}k\Delta t,&\text{if exercise}\\ \tau_{\star},&\text{if continue}\end{cases}$
@@
11:  $v_{0}=\max\left\{\frac{1}{M}\sum_{m=1}^{M}e^{-r\tau_{\star}^{m}}g(\mathbf{S}^{m}_{\tau_{\star}^{m}}),g(\mathbf{S}_{0})\right\}$
@@
$\frac{\partial\phi_{n}^{k}}{\partial w_{j}}(\mathbf{W}^{m}_{k})=\sqrt{\frac{\alpha_{j}}{t_{k}}}\phi_{n^{\prime}}^{k}(\mathbf{W}^{m}_{k})\quad\text{ for }\phi_{n}^{k}=H_{\bm{\alpha}}^{(t_{k})},\phi_{n^{\prime}}^{k}=H_{\bm{\alpha}-\mathbf{e}_{j}}^{(t_{k})},$
@@
Note that $\|I\|_{0}$ can be calculated by $(N_{b}-N_{b,p-1})d$ for maximum order $p$, where $N_{b,p-1}$ represents the number of basis functions with maximum order $p-1$. Figure 1 shows that $\|I\|_{0}$ scales nearly linearly with respect to $N_{b}$. Thus, the complexity of Algorithm 2 is linear in $N_{b}$. Step 8 involves solving a linear system of size $M\times N_{b}$. Using an iterative solver, the cost is $\mathcal{O}(MN_{b})$ if the matrix $A$ is well-conditioned. Step 9 involves matrix-vector multiplication, which has a complexity $\mathcal{O}(MN_{b})$. Hence, with a fixed sample size $M$ and number $N$ of time steps, the total cost of the proposed G-LSM is $\mathcal{O}(NMN_{b})$. We will numerically verify the complexity analysis in Section 6.3.

### v2

@@ 1 Introduction
Nonetheless, the early exercise feature of American options poses significant challenges for computing price and Greeks, especially in high dimensions. One of the most popular methods for high-dimensional American option pricing is the least squares Monte Carlo (LSM) method [18, 28]. Computing Greeks in high dimensions is more involved, and further developments with LSM have been proposed [30, 5]. In this work, building on LSM, we shall develop a simple, fast, and accurate algorithm, termed as gradient-enhanced least squares Monte Carlo (G-LSM) method, cf. Algorithm 1, for computing price and Greeks simultaneously at all time steps for dimensions up to 100. The key methodological innovations includes using sparse Hermite polynomial space with a hyperbolic cross index set as the ansatz space for approximating the continuation value functions (CVFs), and incorporating the gradient information for computing the expansion coefficients.
Next we elaborate the two methodological innovations. First, the main obstacle of using gradient information lies in the computational expense: a $d$-variate price has $d$ partial derivatives, which grows quickly with $d$, especially when the derivative evaluation is costly. With the proposed sparse Hermite polynomial ansatz space, the derivatives of polynomial bases can be obtained at almost no extra cost, cf. Eq. 9 below. This allows greatly reducing the computational cost. Second, although using a polynomial ansatz space for the CVFs as LSM, G-LSM finds the expansion coefficients via solving a linear least squares problem enhanced by the gradient (and hence the name G-LSM), by minimizing the mean squared error between the approximate and exact value functions at $t_{k+1}$, cf. Eq. 15 below. This differs markedly from LSM, which approximates the conditional expectations by projection and minimizes the mean squared error of approximating CVF at $t_{k}$. Numerical experiments show that this choice can achieve better accuracy in price, Greeks, and optimal exercise strategies than LSM. Although other classes of basis functions, e.g., Chebyshev, Legendre, and Laguerre polynomials, might be applicable, we propose to adopt sparse Hermite polynomials as the ansatz space for CVF because (1) CVF can be reformulated as a function of independent Brownian motions with its best approximation error ensured by Lemma 5.1 and (2) the gradient of Hermite polynomials can be evaluated very efficiently, which is the key point to enable the algorithm to work in high dimensions.
@@
The proposed G-LSM scheme is closely linked to the regression-based Monte Carlo method for solving BSDEs [12]. It involves solving a sequence of least squares problems backward in time, where a linear combination of general basis functions is used to approximate each component of the solution $(Y_{t_{k}},Z_{1,t_{k}},\dots,Z_{d,t_{k}})$ to the BSDE. However, this scheme may become infeasible for high dimensions [15]. In sharp contrast, in the $d$-dimensional case, the G-LSM method has a computational complexity $d$ times less that of the approach in [12], while maintains the stability of least squares regression. See Remark 4.7 for further discussions on the complexity and stability. The efficiency of the scheme [12] for high-dimensional problems was enhanced by implementing a neural network in the scheme [15], which has demonstrated impressive empirical performance.
@@
Throughout, bold and plain letters represent multi-variable and scalars, respectively, and the capital and bold letters $\mathbf{S}$, $\tilde{\mathbf{X}}$ and $\mathbf{W}$ denote random vectors. The notation $\mathbf{a}\cdot\mathbf{b}$ denotes the dot product of two vectors $\mathbf{a}$ and $\mathbf{b}$, $\operatorname{Tr}(A)$ the trace of a square matrix $A$, and vector transpose. The notation $\nabla_{\mathbf{x}}f(t,\mathbf{x})$ and $\operatorname{Hess}_{\mathbf{x}}f(t,\mathbf{x})$ denote respectively the gradient and Hessian of $f$ with respect to $\mathbf{x}$. For a multi-index $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{d})^{\top}\in\mathbb{N}_{0}^{d}$, we denote $|\bm{\alpha}|:=\alpha_{1}+\dots+\alpha_{d}$. For a positive-valued integrable function $\omega:\mathbb{R}^{d}\to\mathbb{R}^{+}$, the weighted space $L^{2}_{\omega}(\mathbb{R}^{d})$ is defined by
@@ 2.1 Option pricing and Greeks
Numerically, the price of Bermudan option is used to approximate the American one. The Bermudan option can be exercised at finite discrete times $0=t_{0}<t_{1}<\dots<t_{N}=T$ with $\Delta t:=t_{k+1}-t_{k}$ for all $k=0,1,\dots,N-1$. Using dynamic programming principle or Snell envelope theory [26, Section 1.8.4], the Bermudan price function ${v}_{t_{k}}$ at time $t_{k}$ is given by the following backward induction:
@@
[…]
@@
The options delta and gamma are defined to be the first and second order derivatives of the price function $v_{t_{k}}$ with respect to the price of underlying assets:
(3) $\Delta_{t_{k}}$ $=\nabla v_{t_{k}}(\mathbf{s})=\left(\frac{\partial v_{t_{k}}(\mathbf{s})}{\partial s_{1}},\dots,\frac{\partial v_{t_{k}}(\mathbf{s})}{\partial s_{d}}\right)^{\top}\quad\mbox{and}\quad\Gamma^{ij}_{t_{k}}$ $=\frac{\partial^{2}v_{t_{k}}(\mathbf{s})}{\partial s_{j}\partial s_{i}}.$
We consider all exercise and continuation values of Bermudan option discounted to the time $t=0$. Since the discrete-time approximation for American option pricing is always needed in practical computation, we will make no semantic difference between American and Bermudan options, following [13, p. 133].
@@ 3 Sparse Hermite polynomial expansion and gradient
    Let $\omega_{t}$, $t>0$, be the Gaussian density function defined by $\omega_{t}(\mathbf{y}):=\prod_{j=1}^{d}\rho_{t}(y_{j})$ with $\rho_{t}(y):=\frac{1}{\sqrt{2\pi t}}\exp(-\frac{y^{2}}{2t})$. The CVF $c_{t_{k}}$ satisfies $c_{t_{k}}\in{L}^{2}_{\omega_{t_{k}}}(\mathbb{R}^{d})$, since the payoff $g(\mathbf{S}_{t_{k}})\in L^{2}(\Omega,\mathcal{F}_{t_{k}},\mathbb{Q})$ has a finite second-order moment. Thus, as an orthonormal basis for $L^{2}_{\omega_{t_{k}}}(\mathbb{R}^{d})$, the set of normalized and generalized Hermite polynomials is a natural choice.
@@
    The CVF $c_{t_{k}}$ has the regularity ${H}^{m}_{\omega_{t_{k}}}(\mathbb{R}^{d})$ for any positive integer $m$ by [31, Lemmas 4.2 and 4.3]. The smoothness of the function implies efficient polynomial approximation.
Next, we recall one-dimensional normalized and generalized Hermite polynomials $H_{n}^{(t)}(y)$ and the tensorized $d$-dimensional polynomials $H_{\bm{\alpha}}^{(t)}(\mathbf{y})$, $\bm{\alpha}\in\mathbb{N}_{0}^{d}$. For $d=1$, using the standard probabilist’s Hermite polynomials $H_{n}(x)=(-1)^{n}e^{\frac{x^{2}}{2}}\frac{\mathrm{d}^{n}}{\mathrm{d}x^{n}}e^{-\frac{x^{2}}{2}}$, we define the $n$th-order normalized and generalized Hermite polynomial $H_{n}^{(t)}$ by
@@ 4.1 Gradient-enhanced Least Squares (G-LS)
First we derive a linear backward stochastic differential equation (BSDE) for the CVF $c_{k}$. It is worth noting that the corresponding BSDE does not contain a generator, and the option price is solely a function of Brownian motion rather than the underlying asset. This fact is crucial for developing the G-LSM using sparse Hermite polynomials.
@@
This is a direct consequence of martingale representation theorem [32, Theorem 2.5.2]. For the sake of completeness, we provide a brief proof. By the Feynman-Kac formula, $f(t,\mathbf{w})$ satisfies a $d$-dimensional parabolic PDE subject to a terminal condition
(10) $\left\{\begin{aligned} \frac{\partial f}{\partial t}(t,\mathbf{w})+\frac{1}{2}\operatorname{Tr}\left(\operatorname{Hess}_{\mathbf{w}}f(t,\mathbf{w})\right)&=0,\quad t\in[t_{k},t_{k+1}),\\ f(t_{k+1},\mathbf{w})&=u_{t_{k+1}}(\mathbf{w}).\end{aligned}\right$.
@@
Remark 4.3.
The Feynman-Kac formula in the classical book [23] requires $C^{2}$ regularity of the terminal condition, which is not satisfied by the value function. The regularity requirement can be relaxed by considering the generalized result of nonlinear Feynman-Kac formula in [32, p. 105], which only requires the terminal condition to be uniformly Lipschitz continuous.
Next, we construct an approximation of the CVF ${c}_{k}$ by matching the terminal condition over the interval $t_{k},t_{k+1}]$. By Theorem 4.1, the terminal value over $t_{k},t_{k+1}]$ can be approximated by the Euler discretization:
@@ 4.2 Gradient-enhanced Least Squares Monte Carlo (G-LSM)
where $\{\mathbf{W}^{m}_{k}\}_{m=1}^{M}$ are $M$ independent paths of the Gaussian random process $\mathbf{W}_{k}$ and $\Delta\mathbf{W}^{m}_{k}:=\mathbf{W}^{m}_{k+1}-\mathbf{W}^{m}_{k}$ is the $m$th path of the increment $\Delta\mathbf{W}_{k}$. Let $A_{k}\in\mathbb{R}^{M\times N_{b}}$, $\bm{\beta}_{k}\in\mathbb{R}^{N_{b}}$ and $\mathbf{\hat{u}}_{k+1}\in\mathbb{R}^{M}$ with their components defined by
(16) $(A_{k})_{m,n}$ $=\phi_{n}^{k}(\mathbf{W}^{m}_{k})+\nabla\phi_{n}^{k}(\mathbf{W}^{m}_{k})\cdot\Delta\mathbf{W}^{m}_{k},$
@@
(17) $\bm{\beta}_{k}=\operatorname*{argmin}_{\bm{\beta}}\|A_{k}\bm{\beta}-\mathbf{\hat{u}}_{k+1}\|_{2}^{2}=(A_{k}^{\top}A_{k})^{-1}A_{k}^{\top}\mathbf{\hat{u}}_{k+1}.$
@@
[…]
@@
[…]
@@
(18) $\frac{\partial\phi_{n}^{k}}{\partial w_{j}}(\mathbf{W}^{m}_{k})=\sqrt{\frac{\alpha_{j}}{t_{k}}}\phi_{n^{\prime}}^{k}(\mathbf{W}^{m}_{k})\quad\text{ for }\phi_{n}^{k}=H_{\bm{\alpha}}^{(t_{k})},\phi_{n^{\prime}}^{k}=H_{\bm{\alpha}-\mathbf{e}_{j}}^{(t_{k})},$
@@
Note that $\|I\|_{0}$ can be calculated by $(N_{b}-N_{b,p-1})d$ for maximum order $p$, where $N_{b,p-1}$ represents the number of basis functions with maximum order p1p-1. Figure 1 shows that $\|I\|_{0}$ scales nearly linearly with respect to $N_{b}$. Thus, the complexity of Algorithm 2 is linear in $N_{b}$. Step 8 involves solving a linear system of size $M\times N_{b}$. Using an iterative solver, the cost is $\mathcal{O}(MN_{b})$ if the matrix $A$ is well-conditioned. Step 9 involves matrix-vector multiplication, which has a complexity $\mathcal{O}(MN_{b})$. Hence, with a fixed sample size $M$ and number $N$ of time steps, the total cost of the proposed G-LSM is $\mathcal{O}(NMN_{b})$. We will numerically verify the complexity analysis in Section 6.3.
@@
4.3 Stability of the system of linear equations (17)
Next, we derive the stability of the system of linear equations (17) by estimating the condition number $\mathrm{cond}(A_{k}^{\top}A_{k})$ (under the standard Euclidean norm) of the matrix $A_{k}^{\top}A_{k}$ defined in (16) when the number $M$ of simulation paths is large. This result underpins the efficiency of the standard conjugate gradient method [14] for solving the least-squares problem (the associated normal equation, to be more precise).
Theorem 4.5.
Let $A_{k}\in\mathbb{R}^{M\times N_{b}}$ be defined in (16). Then there holds
$\lim_{M\to\infty}\operatorname{cond}(A_{k}^{\top}A_{k})=1+\frac{p+d-1}{k}\quad\text{almost\; surely.}$
Proof 4.6.
By definition, the $(i,j)$ entry of matrix $A_{k}^{\top}A_{k}$ is given by
(19) $(A_{k}^{\top}A_{k})_{i,j}=\sum_{m=1}^{M}\left(\phi_{i}^{k}(\mathbf{W}_{k}^{m})+\nabla\phi_{i}^{k}(\mathbf{W}_{k}^{m})\cdot\Delta\mathbf{W}_{k}^{m}\right)\left(\phi_{j}^{k}(\mathbf{W}_{k}^{m})+\nabla\phi_{j}^{k}(\mathbf{W}_{k}^{m})\cdot\Delta\mathbf{W}_{k}^{m}\right).$
By the vanilla Monte Carlo approximation, we obtain
$\quad\lim_{M\to\infty}\frac{1}{M}\sum_{m=1}^{M}\left(\phi_{i}^{k}(\mathbf{W}_{k}^{m})+\nabla\phi_{i}^{k}(\mathbf{W}_{k}^{m})\cdot\Delta\mathbf{W}_{k}^{m}\right)\left(\phi_{j}^{k}(\mathbf{W}_{k}^{m})+\nabla\phi_{j}^{k}(\mathbf{W}_{k}^{m})\cdot\Delta\mathbf{W}_{k}^{m}\right)$
$=\mathbb{E}\left[\phi_{i}^{k}(\mathbf{W}_{k})\phi_{j}^{k}(\mathbf{W}_{k})+\nabla\phi_{i}^{k}(\mathbf{W}_{k})\cdot\nabla\phi_{j}^{k}(\mathbf{W}_{k})(\Delta\mathbf{W}_{k})^{2}\right]$
$\quad+\mathbb{E}\left[\left(\phi_{i}^{k}(\mathbf{W}_{k})\nabla\phi_{j}^{k}(\mathbf{W}_{k})+\phi_{j}^{k}(\mathbf{W}_{k})\nabla\phi_{i}^{k}(\mathbf{W}_{k})\right)\cdot\Delta\mathbf{W}_{k}\right]a.s.$
(20) $=\mathbb{E}\left[\phi_{i}^{k}(\mathbf{W}_{k})\phi_{j}^{k}(\mathbf{W}_{k})+\nabla\phi_{i}^{k}(\mathbf{W}_{k})\cdot\nabla\phi_{j}^{k}(\mathbf{W}_{k})(\Delta\mathbf{W}_{k})^{2}\right],$
where the last equality is due to the vanishing mean of Brownian motion increment $\mathbb{E}[\Delta\mathbf{W}_{k}]=0$.
Furthermore, since $\phi_{i}^{k}$ and $\phi_{j}^{k}$ are generalized and normalized Hermite polynomials, then $\mathbb{E}[\phi_{i}^{k}(\mathbf{W}_{k})\phi_{j}^{k}(\mathbf{W}_{k})]=\delta_{ij}$. Then the identity $\mathbb{E}[(\Delta\mathbf{W}_{k})^{2}]=\Delta t$ and (18) lead to
$\mathbb{E}[\nabla\phi_{i}^{k}(\mathbf{W}_{k})\cdot\nabla\phi_{j}^{k}(\mathbf{W}_{k})(\Delta\mathbf{W}_{k})^{2}]=\mathbb{E}\left[\sum_{\ell=1}^{d}\partial_{\ell}\phi_{i}^{k}(\mathbf{W}_{k})\partial_{\ell}\phi_{j}^{k}(\mathbf{W}_{k})\right]\Delta t=0,\quad\text{if }i\neq j.$
This, together with (19) and (20), implies that if $i\neq j$, then $\lim_{M\to\infty}(A_{k}^{\top}A_{k})_{i,j}=0$ a.s., i.e., the off-diagonal entries are nearly vanishing when $M\to\infty$. To estimate the diagonal entries, let $\phi_{i}^{k}(\mathbf{W_{k}})=H_{\bm{\alpha}}^{(t_{k})}$ for some multi-index $\bm{\alpha}$. Using the first order derivative formula (18) and $t_{k}=k\Delta t$, we derive
(21) $\mathbb{E}\left[\sum_{\ell=1}^{d}\partial_{\ell}\phi_{i}^{k}(\mathbf{W}_{k})\partial_{\ell}\phi_{i}^{k}(\mathbf{W}_{k})\right]\Delta t$ $=\Delta t\sum_{\ell=1}^{d}\mathbb{E}\left[\left(\partial_{\ell}H_{\bm{\alpha}}^{(t_{k})}(\mathbf{W}_{k})\right)^{2}\right]$
$=\Delta t\sum_{\ell=1}^{d}\frac{\alpha_{\ell}}{t_{k}}\mathbb{E}\left[\left(H_{\bm{\alpha}-\mathbf{e}_{\ell}}^{(t_{k})}(\mathbf{W}_{k})\right)^{2}\right]=\Delta t\frac{|\bm{\alpha}|}{t_{k}}=\frac{|\bm{\alpha}|}{k}.$
By combining (21) with (20), we obtain an estimate of the diagonal entry,
$\lim_{M\to\infty}\frac{1}{M}(A_{k}^{\top}A_{k})_{i,i}=(1+\frac{|\bm{\alpha}|}{k})\;a.s.$
Here, the involved approximation error is only due to the Monte Carlo method. Thus, the matrix $A_{k}^{\top}A_{k}$ is approximately a diagonal matrix when the sample size $M$ is large. Hence, the definition of condition number implies
$\lim_{M\to\infty}\operatorname{cond}(A_{k}^{\top}A_{k})=\left(1+\frac{\max_{\bm{\alpha\in I}}|\bm{\alpha}|}{k}\right)\Big{/}\left(1+\frac{\min_{\bm{\alpha\in I}}|\bm{\alpha}|}{k}\right)=1+\frac{p+d-1}{k}a.s.$
Here, $I$ is the hyperbolic index set defined in (8). This proves the assertion.
Remark 4.7.
We give a detailed comparison of the proposed G-LSM method with the numerical scheme in [12] for solving BSDEs in terms of computational complexity. The scheme in [12] solves the least squares problem
(22) $\min_{\alpha_{0},\alpha_{1},\dots,\alpha_{d}}\frac{1}{M}\sum_{m=1}^{M}\left(Y_{t_{k+1}}^{N,I,I,M,m}-\alpha_{0}\cdot p_{0,k}^{m}-\sum_{l=1}^{d}\alpha_{l}\cdot p_{l,k}^{m}\Delta W_{l,k}^{m}\right)^{2},$
where each component of the BSDE solution $(Y_{t_{k}},Z_{1,t_{k}},\dots,Z_{d,t_{k}})$ is approximated by the linear combination of distinct basis functions, i.e., $\alpha_{l}\cdot p_{l,k}$. Suppose each approximation uses $N_{b}$ basis functions. Then $\alpha_{l}$ is a vector of length $N_{b}$ for each $l\in\{0,1,\dots,d\}$. In total, the least squares problem (22) involves $(d+1)\times N_{b}$ unknowns for a $d$-dimensional problem. In contrast, the least squares problem in (15) is equivalent to
(23) $\min_{\beta}\frac{1}{M}\sum_{m=1}^{M}\left(\hat{u}_{k+1}(\mathbf{W}_{k+1}^{m})-\beta\cdot(A_{k})_{m,\cdot}\right)^{2},$
where only the CVF is approximated by the linear combination of sparse Hermite basis functions, i.e, the unknown $\beta\cdot\phi^{k}$ with $\beta\in\mathbb{R}^{N_{b}}$ and $\phi^{k}$ being the vector of basis functions. Here, $(A_{k})_{m,\cdot}$ is the $m$-th row of the matrix $A_{k}$ in the linear system, which can be efficiently calculated by Algorithm 2. To the best of our knowledge, this idea is new in literature. We can see that the least squares problem (23) includes $N_{b}$ unknowns for a $d$-dimensional problem. Compared with $(d+1)\times N_{b}$ unknowns in the system (22), the least squares problem (23) incurs much lower computational complexity. Furthermore, the empirical regression matrix for the problem (22) may not be invertible when the sample size becomes large as mentioned in [12, page 2178], while the regression matrix for (23) is well-conditioned for large sample size as proved in Theorem 4.5. In summary, in comparison with the scheme in [12], the proposed G-LSM method has $d$ times less computational complexity for a $d$-dimensional problem, and enjoys guaranteed stability when $M\to\infty$.